# arXiv:2603.03192v1[cs.CV] 03 Mar 2026

### v1

@@ 2.2 Multimodal Policy Optimization
V-DPO [xie2024vdpomitigatinghallucinationlarge] mitigates vision–language hallucinations by injecting video guidance into preference learning through a divergence term that discourages reliance on text-only policies and uses image-contrast pairs to reinforce visual grounding; mDPO [wang-etal-2024-mdpo] strengthens multimodal conditioning via conditional preference terms comparing responses under informative versus degraded images and a reward anchor that prevents down-weighting correct answers; VistaDPO [huang2025_vistadpo] extends preference optimization to videos using hierarchical preferences across instances, time, and regions; and OmniDPO [chen2025omnidpo] generalizes to audio–visual–language settings with text-preference and multimodal-preference pairs that promote attention to visual and auditory evidence. In contrast, our MoD-DPO method is objective-centric and modality-aware, enforcing invariance to perturbations in irrelevant modalities, sensitivity to corruptions in the relevant one, and incorporating a text-only debiasing penalty that directly counters dominant text priors driving multimodal hallucination.
@@ 3 Preliminary
$r(x,y)=\beta\,\frac{\log\pi(y\mid x)}{\log\pi_{\text{ref}}(y\mid x)}+\beta\,Z(x),$ (2)
@@
$\small\mathcal{L}_{\text{DPO}}=-\mathbb{E}\Bigg[\sigma\Bigg(\beta\Bigg(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\Bigg)\Bigg)\Bigg]$ (3)
@@ 4.1 Decoupling Modality Inputs
$\pi^{\star}_{\theta}(y\mid a,v,x^{v})$ $\propto\;\exp\!\big(r(a,v,x^{v},y)\big)\,\pi_{\text{ref}}(y\mid a,v,x^{v})^{\beta/\tau}$ (6)
@@
where $\tau=\beta+\beta_{\text{inv}}+\beta_{\text{sens}}$. Note that this objective is only for prompts related to the visual modality. We can write a similar form for prompts $x^{a}$ related to the audio modality as $\mathcal{L}_{\text{MoD}}^{a}$ (Supp. A.4). The final loss function for decoupling modality inputs is as follows,
@@ 6.2 Ablation Studies
  • Prompt Pool ........................................................................................................................................................................Supp. E
@@ A.1 Finding the Optimal Policy
We can now group the terms by $\log p_{\theta}(y)$ and collect the constants. Additionally, we define $\tau\triangleq\beta+\beta_{\mathrm{inv}}+\beta_{\mathrm{sens}}$ for simplicity. Then Eq. 15 becomes,
@@ D.3 Multi-turn omni-modal results
Appendix E Prompt Pool

### v2

# MoD-DPO: Towards Mitigating Cross-modal Hallucinations in Omni LLMs using Modality Decoupled Preference Optimization

@@ 2.2 Multimodal Policy Optimization
V-DPO [xie2024vdpomitigatinghallucinationlarge] mitigates vision–language hallucinations by injecting video guidance into preference learning through a divergence term that discourages reliance on text-only policies and uses image-contrast pairs to reinforce visual grounding; mDPO [wang-etal-2024-mdpo] strengthens multimodal conditioning via conditional preference terms comparing responses under informative versus degraded images and a reward anchor that prevents down-weighting correct answers; VistaDPO [huang2025_vistadpo] extends preference optimization to videos using hierarchical preferences across instances, time, and regions; and OmniDPO [chen2025omnidpo] generalizes to audio–visual–language settings with text-preference and multimodal-preference pairs that promote attention to visual and auditory evidence. AVEm-DPO [chaubey2026avere] propose to build explicit modality-specific preferences to improve audiovisual emotion understanding. In contrast, our MoD-DPO method is objective-centric and modality-aware, enforcing invariance to perturbations in irrelevant modalities and sensitivity to corruptions in the relevant one.
@@ 3 Preliminary
$r(x,y)=\beta\,\log\frac{\pi(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x),$ (2)
@@
$\small\mathcal{L}_{\text{DPO}}=-\mathbb{E}\Bigg[\log\sigma\Bigg(\beta\Bigg(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\Bigg)\Bigg)\Bigg]$ (3)
@@ 4.1 Decoupling Modality Inputs
$\pi^{\star}_{\theta}(y\mid a,v,x^{v})$ $\propto\;\exp\!\big(r(a,v,x^{v},y)/\tau\big)\,\pi_{\text{ref}}(y\mid a,v,x^{v})^{\beta/\tau}$ (6)
@@
where $\tau=\beta+\beta_{\text{inv}}-\beta_{\text{sens}}$. Note that this objective is only for prompts related to the visual modality. We can write a similar form for prompts $x^{a}$ related to the audio modality as $\mathcal{L}_{\text{MoD}}^{a}$ (Supp. A.4). The final loss function for decoupling modality inputs is as follows,
@@ 6.2 Ablation Studies
  • Discussion on Cross-Modal Synergy ........................................................................................................................................................................Supp. E
  • Prompt Pool ........................................................................................................................................................................Supp. F
@@ A.1 Finding the Optimal Policy
We can now group the terms by $\log p_{\theta}(y)$ and collect the constants. Additionally, we define $\tau\triangleq\beta+\beta_{\mathrm{inv}}-\beta_{\mathrm{sens}}$ for simplicity. Then Eq. 15 becomes,
@@ D.3 Multi-turn omni-modal results
Appendix E Discussion on Cross-Modal Synergy
The MoD-DPO objective defined in Sec. 4.1 mitigates spurious inter-modality correlations by introducing explicit regularization terms that enforce modality invariance and sensitivity. However, many multimodal tasks benefit from constructive synergy between audio and visual modalities for richer audiovisual understanding and reasoning. Extending the MoD-DPO framework with mechanisms to selectively promote beneficial cross-modal interactions, while still suppressing spurious ones, remains an important direction for future work.
Appendix F Prompt Pool